OpenStudy (anonymous):

3x^3+5x^2-3x-5=0 factor by grouping.

OpenStudy (anonymous):

\[x^2(3x+5)-1(3x+5)=0\] \[(3x+5)(x^2-1)=0\] \[3x+5=0,x^2-1=0\] \[x=\frac{-5}{3},x=1\]

OpenStudy (anonymous):

\[x^2-1=0\] \[x=1,x=-1\]

OpenStudy (anonymous):

sorry i forgot \[\pm\]

OpenStudy (anonymous):

3x^3+5x^2-3x-5=0 3x^3-3x+5x^2-5=0 3x(x^2-1)+5(x^2-1)=0 (3x+5)(x^2-1)=0 (3x+5)(x+1)(x-1)=0